# Two critical temperatures conundrum in ${\rm La}^{\vphantom{\dagger}}_{1.83}{\rm Sr}^{\vphantom{\dagger}}_{0.17}{\rm CuO}^{\vphantom{\dagger}}_{4}$

### v5

@@ 1 Introduction
A homogeneous three-dimensional superconductor is expected to exhibit a single transition temperature ${T_{\rm c}}$ at which the order parameter, $\Delta(T)$, and all the superconducting stiffness components vanish [1, 2]. In this regard, recent measurements of the $ab$-plane ($\rho^{\parallel}$) and $c$-axis ($\rho^{\perp}$) stiffnesses of La1.875Sr0.125CuO4 (LSCO) crystals by Kapon et. al. [3] have been puzzling. Counter to the expectation above, $\rho^{\perp}$ was seen to vanish at ${T_{\rm c}}^{\perp}$, which is about $0.64$ K below the vanishing temperature ${T_{\rm c}}^{\parallel}$ of $\rho^{\parallel}$.
@@
This paper explores finite size effects experimentally and theoretically. We report systematic stiffness measurements near ${T_{\rm c}}$ for LSCO rings with widths $L_{a},L_{c}$ ranging between $L=0.1$ to 1 millimeter. ${T_{\rm c}}^{\parallel}$ is found to be weakly dependent on $L_{c},L_{a}$. In contrast, a significant reduction of ${T_{\rm c}}^{\perp}$ for decreasing width $L_{a}$ is observed. This behavior is not expected for layer-correlated inhomogeneity. The relatively strong finite size effect demands theoretical explanation.
@@
We compare our theoretical analysis to the experimental values ${T_{\rm c}}^{\perp}(L_{a})$, and use the fit to estimate of the anisotropy parameter of LSCO near its thermodynamic ${T_{\rm c}}$.
@@ 2 Experiments
La2-xSrxCuO4 is known to grow in large single crystals allowing significant size reductions. A bunch of crystals with different doping where grown using an image furnace, oriented with a Laue camera, and cut into rings with a femtosecond laser cutter. For each doping two rings, labelled by $a$ and $c$, were prepared with their coil axes parallel ($a$) and perpendicular ($c$) to the CuO2 planes. The width and height of the rings was varied by the laser, or polished down to the geometries shown in Fig. 2a,b.
Figure 2: Experimental configuration and normalized magnetizations of LSCO rings for a fixed 1 mA current in the coil. (a) The interior of the $a$-ring. The CuO2 planes are parallel to the ring’s symmetry axis. This ring is sequentially polished to reduce the layers’ width $L_{a}$ in the bottleneck region. (b) The interior of the $c$-ring. The CuO2 planes are perpendicular to the ring’s symmetry axis. This ring is polished along two planes which varies both $L_{a}$ and $L_{c}$. (c) Photograph of an $a$-ring with two cut planes. $L_{c}^{\rm eff}$ defines the effective aspect ratio in the bottleneck region. (d) Magnetization $m^{a}$ of $a$-rings with variable $L_{a}$. The apparent stiffness vanishing temperatures are denoted by ${\widetilde{T}}_{\rm c}^{\perp}(L_{a})$. (e) Magnetization $m^{c}$ of $c$-rings. Except for the narrowest sample $L_{a}\!=\!0.09$ (which is suspected of containing a traversing cut), the magnetizations near their transition are insensitive to $L_{a}$. The tail of width $(\delta{T}_{\rm c}^{\parallel})_{\rm dis}\simeq 0.5$ K is assumed to reflect some layer-correlated disorder, which is a smaller effect than the finite size dependence of the $a$-rings’ ${\widetilde{T}}_{\rm c}^{\perp}$. $\bar{T}_{\rm c}^{\parallel}$ is averaged in-plane transition temperature (see Section 6 and Appendix 7).
@@ 3 Layered 3dXY model
As mentioned before, we model the phase fluctuations of LSCO near ${T_{\rm c}}$ by the classical 3dXY Hamiltonian on a tetragonal lattice,
@@ 6 Comparison of Theory to Experiments
Figure 6: Comparison of experimental results for LSCO rings of Fig. 2 and theoretical prediction of Eqs. (LABEL:deltaTc) and (19). Crosses: The apparent $c$-axis transition temperature shifts ${\widetilde{T}}_{\rm c}^{\perp}$ of the $a$-rings, as determined in Fig. 2(d). $L_{a}$ are the bottleneck widths of the $ab$ planes. Line: the least square fit using $\alpha^{\rm fit}=4.1\times 10^{-5}$. The offset of the reduced temperature $0.015(4)$ agrees with the estimated layer-correlated disorder (see text). We use the mean field exponent $\beta={1\over 2}$, due to the narrow Ginzburg critical regime near ${T_{\rm c}}$.
For the experimental LSCO crystals, the millimeter width corresponds to $\sim 10^{5}$ effective lattice constants, and the anistropy parameter will turn out to be $\alpha<10^{-4}$, which yields an unobservable narrow Ginsburg critical region. Hence we shall fit the Eq. LABEL:deltaTc with the mean field exponent $\beta=0.5$.
@@ 7 Discussion and Summary
We note that $\alpha^{\rm fit}$ parametrizes the effective Hamiltonian near ${T_{\rm c}}$. We compare it to the zero temperature anisotropy parameter reported for optimally doped LSCO (for Sr0.15) in Ref. [29],

### v6

@@ 1 Introduction
A homogeneous three-dimensional superconductor is expected to exhibit a single transition temperature ${T_{\rm c}}$ at which the order parameter, $\Delta(T)$, and all the superconducting stiffness components vanish [1, 2]. In this regard, recent measurements of the $ab$-plane ($\rho^{\parallel}$) and $c$-axis ($\rho^{\perp}$) stiffnesses of La1.875Sr0.125CuO4 crystals by Kapon et. al. [3] have been puzzling. Counter to the expectation above, $\rho^{\perp}$ was seen to vanish at ${T_{\rm c}}^{\perp}$, which is about $0.64$ K below the vanishing temperature ${T_{\rm c}}^{\parallel}$ of $\rho^{\parallel}$.
@@
This paper explores finite size effects experimentally and theoretically. We report systematic stiffness measurements near ${T_{\rm c}}$ for La1.83Sr0.17CuO4 rings with widths $L_{a},L_{c}$ ranging between $L=0.1$ to 1 millimeter. ${T_{\rm c}}^{\parallel}$ is found to be weakly dependent on $L_{c},L_{a}$. In contrast, a significant reduction of ${T_{\rm c}}^{\perp}$ for decreasing width $L_{a}$ is observed. This behavior is not expected for layer-correlated inhomogeneity. The relatively strong finite size effect demands theoretical explanation.
@@
We compare our theoretical analysis to the experimental values ${T_{\rm c}}^{\perp}(L_{a})$, and use the fit to estimate of the anisotropy parameter of La1.83Sr0.17CuO4 near its thermodynamic ${T_{\rm c}}$.
@@ 2 Experiments
La2-xSrxCuO4 is known to grow in large single crystals allowing significant size reductions. Therefore, powder of different doping is prepared from stoichiometric ratios of 99.99% pure CuO, La2O3, and SrCO3 to make feed and seed rods. This powder is turned into a single crystal using an image furnace with four elliptic mirrors focusing 300 W halogen lamps. The growth was stabilised over 100 hr without any change of the lamp 59% power. Growth rate of 1.0 mm/h, down-ward translation of 0.15 mm/h, and rotation in opposite directions at 15 rmp were used. The emerging crystals looked like Fig. 1 of Ref. [15]. After the growth, the crystals were annealed in argon environment at $T=850$ C for 120 hr to release internal stress. Finally, the crystals were oriented with a Laue camera, and cut into rings with a femtosecond laser cutter. For each doping two rings, labelled by $a$ and $c$, were prepared with their coil axes parallel ($a$) and perpendicular ($c$) to the CuO2 planes. The width and height of the rings was varied by the laser, or polished down to the geometries shown in Fig. 2a,b.
Figure 2: Experimental configuration and normalized magnetizations of La1.83Sr0.17CuO4 rings for a fixed 1 mA current in the coil. (a) The interior of the $a$-ring. The CuO2 planes are parallel to the ring’s symmetry axis. This ring is sequentially polished to reduce the layers’ width $L_{a}$ in the bottleneck region. (b) The interior of the $c$-ring. The CuO2 planes are perpendicular to the ring’s symmetry axis. This ring is polished along two planes which varies both $L_{a}$ and $L_{c}$. (c) Photograph of an $a$-ring with two cut planes. $L_{c}^{\rm eff}$ defines the effective aspect ratio in the bottleneck region. (d) Magnetization $m^{a}$ of $a$-rings with variable $L_{a}$. The apparent stiffness vanishing temperatures are denoted by ${\widetilde{T}}_{\rm c}^{\perp}(L_{a})$. (e) Magnetization $m^{c}$ of $c$-rings. Except for the narrowest sample $L_{a}\!=\!0.09$ (which is suspected of containing a traversing cut), the magnetizations near their transition are insensitive to $L_{a}$. The tail of width $(\delta{T}_{\rm c}^{\parallel})_{\rm dis}\simeq 0.5$ K is assumed to reflect some layer-correlated disorder, which is a smaller effect than the finite size dependence of the $a$-rings’ ${\widetilde{T}}_{\rm c}^{\perp}$. $\bar{T}_{\rm c}^{\parallel}$ is averaged in-plane transition temperature (see Section 6 and Appendix 7).
@@ 3 Layered 3dXY model
As mentioned before, we model the phase fluctuations of La1.83Sr0.17CuO4 near ${T_{\rm c}}$ by the classical 3dXY Hamiltonian on a tetragonal lattice,
@@ 6 Comparison of Theory to Experiments
Figure 6: Comparison of experimental results for La1.83Sr0.17CuO4 rings of Fig. 2 and theoretical prediction of Eqs. (LABEL:deltaTc) and (19). Crosses: The apparent $c$-axis transition temperature shifts ${\widetilde{T}}_{\rm c}^{\perp}$ of the $a$-rings, as determined in Fig. 2(d). $L_{a}$ are the bottleneck widths of the $ab$ planes. Line: the least square fit using $\alpha^{\rm fit}=4.1\times 10^{-5}$. The offset of the reduced temperature $0.015(4)$ agrees with the estimated layer-correlated disorder (see text). We use the mean field exponent $\beta={1\over 2}$, due to the narrow Ginzburg critical regime near ${T_{\rm c}}$.
For the experimental La1.83Sr0.17CuO4 crystals, the millimeter width corresponds to $\sim 10^{5}$ effective lattice constants, and the anistropy parameter will turn out to be $\alpha<10^{-4}$, which yields an unobservable narrow Ginsburg critical region. Hence we shall fit the Eq. LABEL:deltaTc with the mean field exponent $\beta=0.5$.
@@ 7 Discussion and Summary
We note that $\alpha^{\rm fit}$ parametrizes the effective Hamiltonian near ${T_{\rm c}}$. We compare it to the zero temperature anisotropy parameter reported for optimally doped La2-xSrxCuO4 (for Sr0.15) in Ref. [30],